# Quantum and classical algorithms for daily railway rolling stock circulation plans

### v1

@@ Abstract
We study daily rolling stock circulation planning for electric multiple units (EMUs) on a regional passenger network, focusing on services where identical EMUs may be coupled in pairs on selected routes. Motivated by the operational needs of the regional operator Silesian Railways in Poland, we formulate an acyclic mixed‑integer linear program on a one‑day horizon that incorporates depot balance constraints, demand‑driven seat and bicycle capacity limits (which is a new aspect requested by the regional operator and local society of passengers), and simple crew availability constraints. The model is designed to support both baseline planning and disruption management under increased passenger demand. Using a graph–hypergraph representation of trips and single or coupled EMU movements, we first solve the problem with a classical ILP solver. We then derive a Quadratic Unconstrained Binary Optimization (QUBO) reformulation - which is frequently used as the input for quantum optimization - and evaluate its solution by quantum annealing on D‑Wave Advantage systems and by the classical quantum‑inspired VeloxQ solver. Computational experiments on real‑world instances from the Silesian network, with up to 404 train trips and 11 EMU types, show that the ILP approach can obtain high‑quality daily circulation plans within at most about 40 minutes, whereas current quantum and quantum‑inspired solvers are restricted to substantially smaller sub‑instances (up to 51 and 78 train trips, respectively) due to embedding and QUBO size limitations. These results quantify the present frontier of QUBO‑based methods for rolling stock circulation and point towards hybrid decision‑support architectures in which quantum or quantum‑inspired optimizers address only local subproblems within a broader classical planning framework.
@@ 1 Introduction
Optimization of processes in railway transport is essential to ensure its efficient operation. The best way to develop such models is to work closely with stakeholders. Therefore, the research problem addressed in this work is based on the specific requirements of the local community of passengers as well as the regional passenger rail operator — Silesian Railways (pl. Koleje Śląskie).
Our goal is to develop a flexible model, yielding a reactive rolling stock plan on an operational time horizon, subject to crew availability constraints, that can respond efficiently to unexpected events. Our model supports different optimization objectives, aiming to balance two primary goals: maximizing the rolling stock reserve and minimizing the operational costs associated with rolling stock usage. Due to the requirements of the local society of passengers and the railway operator, we introduce the bicycle capacity constraint, as the operator serves many tourist destinations.
Rolling stock circulation models are essentially modifications of set-flow models, which do not scale well to large problem instances. In direct implementations of quantum and quantum-inspired approaches, the optimization problem (a variant of the set-flow problem) is formulated as a Quadratic Unconstrained Binary Optimization (QUBO) problem, which also incurs significant overhead for large instances. Recall that the QUBO formulation is the most recognisable input for quantum (quantum-inspired) optimisation; however, we recognise that there are also other approaches, for example, variational quantum algorithms with qubit-efficient encodings [1]. Based on these observations, we assess the current frontier of QUBO-based quantum and quantum-inspired approaches in comparison with classical approaches. Figure (1) describes the workflow of our work.
@@ 1.1 Problem Description
We start the problem description from the practical point of view. The railway operator’s staff, equipped with knowledge of the timetable, the available railway vehicles and their technical and operational parameters, the pool of drivers based in depots, and expected passenger demand, prepare detailed planning tables. These tables assign individual EMU (or pairs of EMUs in some predefined cases) to specific trains (operating on designated routes at specified times) while adhering to all internal and external constraints. Additionally, given the baseline plan, the operator must respond to disruptions, such as increased passenger demand, rolling stock breakdowns, and other issues, in order to minimize the propagation of disturbances throughout the network. We focus on the acyclic rolling stock planning on the operational time horizon with the drivers’ ability aspect, yielding either the original vehicle circulation plan or its version given some disturbances (i.e. increased passenger or bicycle demand). Hereby, we recognise that we do not create an entire driver roster (including driving times, breaks, and many other aspects) but only introduce some constraints to check the drivers’ ability.
A crucial aspect of rolling stock circulation planning is the need for coordination with infrastructure management. The operator values having multiple solutions to the problem, each with different objective values, available within the decision support system (DSS). This allows the dispatcher to make the final decision while maintaining communication with the infrastructure manager.
@@ 1.2 Literature review
In summary, the literature on rolling stock circulation offers a wide range of modelling paradigms, including network-flow formulations, mixed-integer models with flexible train compositions, decomposition-based approaches, and large-scale exact and metaheuristic methods. While many studies address capacity constraints, passenger demand, or operational feasibility, only a limited subset simultaneously considers flexible EMU compositions together with demand-driven capacity requirements in a way that remains computationally tractable for real-world instances. (There is also scarce literature on constraints tied to bicycle capacity.) Moreover, existing works differ substantially in modelling detail, scope, and solution methodology, making it difficult to directly compare contemporary approaches or to assess their frontier for large-scale practical applications. These observations naturally lead to several research gaps, which we discuss in the following section.
@@ 1.3 Research gaps
We recognize the following research gaps. Existing optimization models can grow to sizes that are difficult to handle for realistic instances. In theory, quantum computing promises to address such large-scale problems in a reasonable time once the technology matures. However, we are currently in the NISQ era, in which quantum devices are small and prone to errors. It is therefore important to assess the present stage of development of quantum approaches for the practical problem of railway rolling stock circulation planning. There is little literature on quantum computing or quantum-inspired approaches for the rolling stock circulation planning, to our knowledge, limited to the article [19], where the problem of rolling stock circulation planning on the German Railways network was examined.
As quantum devices are small and prone to error, a hybrid quantum-classical approach [20] has been applied therein and to a somewhat similar railway rescheduling problem [21]. However, such hybrid solvers are proprietary, and it is difficult to assess the potential impact of the quantum or quantum-inspired part of the algorithm on the overall optimization performance.
Concluding, although quantum computing promises to solve large NP-hard problems, there is a lack of research on the actual frontier between classical, quantum-inspired and quantum approaches for railway rolling stock planning. To fill these research gaps, we investigate this frontier (in the context of QUBO formulation) for a practical railway rolling stock planning problem. Our contributions are:
@@
To be complete, as stakeholder-driven research, we enrich the model with the bicycles’ capacity requirements, as the regional operator, beyond commuters, also serves an increasing number of tourists. To place our paper in the literature precisely, please refer to Tab. 1
@@ 2.2 QUBO encoding
As discussed before, we have selected the QUBO-based approach as it is native to the quantum annealing, which, at the current state-of-the-art, can handle relatively large (compared with other paradigms of quantum computing) optimization problems in terms of the number of binary variables. For quantum and quantum-inspired approaches, we encode the problem as the Quadratic Unconstrained Binary Optimization (QUBO). In such implementation, we treat all constraints as soft with large penalties and then post‑filter infeasible solutions, rather than trying to guarantee feasibility by construction. We follow a standard ILP–to–QUBO transformation as in [27].
@@
Using the QUBO encoding, the overhead in the number of variables (compared with the ILP scaling in Eq. (13)) is entirely due to the slack variables, since all ILP decision variables are already binary:
@@
Concluding, the number of QUBO terms scales in the worst case as $|R|^{3}$ and $|T|^{5}$, which is problematic when the number of rolling stock types is large. Hence, it is worth examining the frontier of problem sizes that can still be effectively transformed to QUBO, in order to design future hybrid approaches.
@@ 3 Solution of the actual problem
  • a QUBO encoding, native for quantum annealers and various quantum-inspired techniques.
To demonstrate the frontier of problem sizes, we solve QUBO using both quantum annealing and quantum-inspired approaches.
Quantum annealing [28, 29] is a heuristic algorithm that runs in the framework of adiabatic quantum computing (AQC) [30], targeting optimization problems. In AQC, a system starting in the lowest energy state (ground state) of some initial Hamiltonian $H_{0}$ (a mathematical operator that describes the system’s energy) is likely to remain in the ground state throughout the evolution, provided that the system is evolved slowly enough. Hence, if a problem Hamiltonian $H_{P}$ is introduced gradually, the system is expected to end up in the ground state of $H_{P}$ at the end of the evolution time $T$. Mathematically, the evolution of the system is often described by the time-dependent Hamiltonian
$H(t)=\left(1-\frac{t}{T}\right)H_{0}+\frac{t}{T}H_{P}.$ (28)
or, more generally:
$H(t)=f\left(\frac{t}{T}\right)H_{0}+f\left(\frac{t}{T}\right)H_{P}.$ (29)
where $f$ is the custom “protocol” function, selected in our case to be a linear function.
The quantum annealers provided by D-Wave implement a problem Hamiltonian whose energy is expressed using a 2-local Ising model [31]:
$H_{P}=\sum_{i,j\in\mathcal{E}}J_{ij}\sigma^{z}_{i}\sigma^{z}_{j}+\sum_{i}h_{i}\sigma^{z}_{i},$ (30)
where $\sigma^{z}$ is a Pauli-$Z$ operator acting on the $i$-th qubit, and $J_{ij}$ and $h_{i}$ are real values corresponding to pairwise couplings and the external magnetic field, respectively. The set $\mathcal{E}$ represents the processor graph, i.e., the connectivity graph of a particular quantum device. However, the connectivity of quantum hardware often differs from the connectivity required by the optimization problem. To handle this mismatch, a minor embedding is required to translate the problem topology onto the processor topology [32]. In the D-Wave implementation, one chooses $H_{0}$ to consist of a transverse magnetic field:
$H_{0}=h_{0}\sum_{i}\sigma^{x}_{i}$ (31)
where $\sigma^{x}_{i}$ is the Pauli-$X$ operator acting on the $i$-th qubit.
To see why the QUBO representation in Eqs. (16) - (21) can be implemented straightforwardly on a quantum annealer, observe that a binary variable $y_{i}$ can be easily transformed into a spin variable $s_{i}\in\{-1,1\}$ via $y_{i}=\frac{s_{i}+1}{2}$  [31]. Substituting this into Eq. (14) yields the following optimization problem:
$\min_{\vec{s}}\left(\sum_{i<j}J_{i,j}s_{i}s_{j}+\sum_{i}h_{i}s_{i}\right).$ (32)
Following [33], such a formulation is supposed to be minimized if corresponding to the minimal energy state of the quantum Hamiltonian in Eq. (30).
@@
A potential alternative direction in solving QUBOs is the application of quantum-inspired algorithms [34]. Following this path, we apply the new VeloxQ solver [22], which appears to be effective for large, sparse QUBOs (with a limited number of connections) [25, 24, 26, 23]. This heuristic solver takes a QUBO or Ising instance and returns a series of solutions intended to originate from the problem’s low-energy spectrum (hopefully including at least one feasible solution). VeloxQ is a physics-inspired classical solver for QUBO/Ising problems that mimics energy-minimization dynamics without requiring quantum hardware and achieve high scalability on standard CPUs/GPUs. Compared to D-Wave quantum annealing and simulated bifurcation, it delivers comparable or superior solution quality and runtime, especially for large, arbitrarily connected problem instances.
@@ 3.1 Toy model
The toy model is defined by the timetable $T=\{\tau_{1},\tau_{2},\tau_{3}\}$ where $\tau_{1}$ and $\tau_{2}$ are trains that go $A\rightarrow B$, and $\tau_{3}$ goes $B\rightarrow A$. The timetable is arranged so that any rolling stock that arrives at $B$ can either remain there or return to $A$. Two EMU types $R=\{r_{1},r_{2}\}$ are available at the depot at station $A$ throughout the time horizon, with fleet sizes $N(v_{d})_{r_{1}}=2,N(v_{d})_{r_{2}}=1$ with the seats capacity $p_{1}=70$, $p_{2}=110$. There is also an optional service train $v_{4}$ going from $B$ to $A$ (which can be served only by type $r_{1}$). We assume that $A=2$ drivers are available throughout the horizon; they can drive any of the two EMUs and take any of the trips. Finally, we do not consider bicycles in this example.
@@ 3.2 Practical problems
Results obtained with the quantum-inspired VeloxQ solver [22] are presented in Fig. 9. We observe better performance than D-Wave for small instances; however, large instances were not tractable with this approach. Although the VeloxQ achieves better results than D-Wave, the performance gap remains moderate when compared with the classical ILP approach.
Figure 9: Results of the VeloxQ experiments [22], computational time on VeloxQ solver, all results were feasible and optimal, we used the default setting for instances (toy, 1, 1a, 2, 2a). The computations were performed on 13th Gen Intel(R) Core(TM) i5-1340P @4 GHz, 12 cores, 16 threads.
@@ 4 Discussion
A key requirement from practice is not only to obtain one feasible plan, but to have multiple feasible alternatives available quickly so that dispatchers can account for non-modelled considerations and coordinate with infrastructure management. The toy example illustrates that the model naturally admits multiple feasible “excited” solutions beyond the optimum. This feature aligns well with the behavior of annealing-based approaches, which aim to sample from low-energy regions rather than output a single deterministic solution; in principle, this could be operationally valuable when a DSS needs to offer a small portfolio of near-optimal circulation options (e.g., different reserve usage levels, different coupling choices, or different depot return patterns) under time pressure. However, the experiments also show that obtaining a useful portfolio of feasible solutions from QUBO samplers requires careful constraint encoding (penalty design), feasibility checks, and likely post-processing—otherwise low-energy samples can still violate critical constraints (e.g., slack-variable feasibility).

### v2

# EMU circulation planning for Silesian Railways: case study and a quantum approach

@@ Abstract
We study daily rolling stock circulation planning for electric multiple units (EMUs) on a regional passenger network, focusing on services where identical EMUs may be coupled in pairs on selected routes. Motivated by the operational needs of the regional operator Silesian Railways in Poland, we formulate an acyclic mixed‑integer linear program on a one‑day horizon that incorporates depot balance constraints, demand‑driven seat and bicycle capacity limits (which is a new aspect requested by the regional operator and local society of passengers), and simple crew availability constraints. Using a graph–hypergraph representation of trips and single or coupled EMU movements, we first solve the problem with a classical ILP solver. We then derive a Quadratic Unconstrained Binary Optimization (QUBO) reformulation - which is frequently used as the input for quantum optimization - and evaluate its solution by quantum annealing on D‑Wave Advantage systems and by the classical quantum‑inspired VeloxQ solver. Computational experiments on real‑world instances from the Silesian network, with up to 404 train trips and 11 EMU types, show that the ILP approach can obtain high‑quality daily circulation plans within at most about 40 minutes, whereas current quantum and quantum‑inspired solvers are restricted to substantially smaller sub‑instances (up to 51 and 78 train trips, respectively) due to the large number of terms in the QUBO, and embedding limitations in the case of quantum hardware. These results quantify the present frontier of QUBO‑based methods for rolling stock circulation and point towards hybrid decision‑support architectures in which quantum or quantum‑inspired optimizers address only local subproblems within a broader classical planning framework.
@@ 1 Introduction
Rolling stock circulation planning is an essential part of the planning processes of passenger railway transportation [1]. Given a passenger timetable and rolling stock characteristics, the goal is to optimally assign vehicles to the passenger trains. The so-arising rolling stock circulation plan should obey technological and capacity constraints. The optimization can typically aim at seat capacities, minimization of empty runs, etc. The problem has many variants with a broad literature, which will be reviewed in Section 1.2.
The research problem addressed in this work is based on the specific requirements the regional passenger rail operator — Silesian Railways (pl. Koleje Śląskie), which incorporates results of various forms of feedback from passengers. The company’s traffic is served with EMUs, and the units return to their depots typically on a daily basis. The depots thus typically have a reserve of units. Constraints on the availability and suitability of crews (i.e. train drivers) have also be taken into account. Our goal is to develop a flexible model, yielding a reactive rolling stock plan on an operational time horizon of a few days. Our model supports different optimization objectives, aiming to balance two primary goals: maximizing the rolling stock reserve at the depots and minimizing the operational costs associated with rolling stock usage. As a new element of optimisation aspect, we introduce bicycle capacity constraint. As it will be clear from the literature review, in spite of the broad coverage of passenger railway rolling stock circulaiton optimization, the variant of the problem discussed here, up to our knowledge, has not been dealt with so far.
Rolling stock circulation models are essentially modifications of set-flow models, which do not scale well with the size of problem instances. In spite of that, in the present contribution we could obtain practically useful results using a state-of-the-art solver for constrained integer programs. Meanwhile, as an experimental approach, we have transcribed our model to a Quadratic Unconstrained Binary Optimization (QUBO) form using penalties and slack variables, as it is often done with constrained integer programs [2]. We have solved the so-arising QUBO problems on a hardware quantum annealer and also with a solver implementing advanced physics-based heuristics. On this basis, we assess the current frontier of QUBO-based quantum and quantum-inspired approaches in comparison with the classical approach.
This paper is organized as follows: in the rest of the Introduction, we decribe the problem in more detail and give a brief review of the literature, restricted to works which show the most of resmblance to our problem. In Section 2 introduce the model, whereas Section 3 discusses solutions to actual problem instances. In 4 the results are discussed and conclusions are drawn. Figure (1) describes the workflow of our work.
@@ 1.1 Problem Description
We start the problem description from the practical point of view. The railway operator’s staff, equipped with knowledge of the timetable, the available railway vehicles and their technical and operational parameters, the pool of drivers based in depots, and expected passenger demand, prepare detailed planning tables. These tables assign individual EMU (or pairs of EMUs in some predefined cases) to specific trains (operating on designated routes at specified times) while adhering to all internal and external constraints. We focus on the acyclic rolling stock planning on the operational time horizon with the drivers’ ability aspect, resulting in a vehicle circulation plan. The model can, under certain circumstances, also be used to create an alternative plan for modified circumstances, such as, e.g. increased passenger or bicycle demand, or disturbances. Concerning crews, the model is limited to take into account the number of suitable crews, it is not for crew rostering and scheduling. It only ensures that the appropriate number and type of crews is assigned to the trains and the crew rostering and scheduling will likely to be feasible with the given assingnment.
@@ 1.2 Literature review
In summary, the literature on rolling stock circulation offers a wide range of modelling paradigms, including network-flow formulations, mixed-integer models with flexible train compositions, decomposition-based approaches, and large-scale exact and metaheuristic methods. While many studies address capacity constraints, passenger demand, or operational feasibility, only a limited subset simultaneously considers flexible EMU compositions together with demand-driven capacity requirements in a way that remains computationally tractable for real-world instances. Bike capacity constraints appear to be a new element in the present contribution. Moreover, existing works differ substantially in modelling detail, scope, and solution methodology, making it difficult to directly compare contemporary approaches or to assess their frontier for large-scale practical applications. These observations naturally lead to several research gaps, which we discuss in the following section.
@@ 1.3 Research gaps
We recognize the following research gaps. As of the railway aspect, as already stated, there is no model available in the literature with the exact inputs and outputs required by the actual operator. While our model is conceptually within the range of typical rolling stock circulation models, it has several specific details. Notably, it can be solved for practical instances without column generation techniques.
It is widely recognized in the railway research community that neural network approaches introduce new paradigms in railway optimization, also in modeling approaches. It is also expected that quantum methods will produce a similar paradigm shifts. There is little literature on quantum computing or quantum-inspired approaches for the rolling stock circulation planning, to our knowledge, limited to the article [20], where the problem of rolling stock circulation planning on the German Railways network was examined. That problem, however, is significantly different with respect to input, output, and constraints. Their methodology is less typical in the rolling stock circulation literature, and thus their comparison of classical and quantum approaches does not cover the type of models studied here.
As quantum devices are small and prone to error, a hybrid quantum-classical approach [21] has been applied therein and to a somewhat similar railway rescheduling problem [22]. However, such hybrid solvers are proprietary, and it is difficult to assess the potential impact of the quantum or quantum-inspired part of the algorithm on the overall optimization performance. The present contribution makes a comparison of a pragmatic classical approach with a quantum-compatible one, not only with quantum hardware but also with phyiscs-based heurisitcs.
Our contributions are:
@@
Table 1 provides a brief summary of the similarities and differences of the present contribution to the ones in the literutere most similar to it.
@@ 2.2 QUBO encoding
As discussed before, we have selected the QUBO-based approach as it is native to quantum annealing hardware, which, at the current state of the art, can handle relatively large optimization problems. Hence, for quantum and quantum-inspired approaches, we encode our problem as Quadratic Unconstrained Binary Optimization (QUBO). In such implementation, we treat all constraints as soft with large penalties and then post‑filter infeasible solutions, rather than trying to guarantee feasibility by construction. We follow a standard ILP–to–QUBO transformation as in [2].
@@
Using the QUBO encoding, the overhead in the number of terms (compared with the ILP scaling in Eq. (13)) is entirely due to the slack variables, since all ILP decision variables are already binary:
@@
Concluding, the number of QUBO terms scales in the worst case as $|R|^{3}$ and $|T|^{5}$. This is to be compared with the scaling of the linear formulation in Eq. (11): there the scaling in rolling stock types is linear. As in practical instances $|T|$ is order of 10, this suggests that the QUBO approach will be less practical in case of large instances. Yet it is conceptually interesting as it is one of the native input formats of quantum devices.
@@ 3 Solution of the actual problem
  • a QUBO encoding, as it is native for quantum annealers and various quantum-inspired techniques.
In order to solve ILPs, we opt for the SCIP [28, 29] as it is probably the most efficient non-commercial linear integer programming solver at the moment. According to our experience, this solver could be used even in production in the studied railway environment.
As for QUBOs, they have a relevant literature also in classical optimization (see e.g. the monograph edited by Punnen [30]), and they can also be solved with classical solvers. The aforementioned overhead in the number of variables of the QUBO model suggests that our direct reformulation of the ILP to QUBO will not be practical for bigger instances. Indeed, our aim is to explore the behavior of quantum hardware and physics-based QUBO heuristics rather than to consider this as a practical solution for big instances.
Quantum annelaers, such as DWave which is used in the present contribution expoloit the equivalence of QUBOs with Ising models prevalently known in physics. A quantum annealer actually realizes a configurable pyhsical setup whose lowest-energy state gives an optimal configuration for QUBOs[31]. A quantum annealer is not a universal solver of QUBOs, however, in the case of certain instances, it can potentially be more efficient than classical algorithms. It is heuristic by nature: it provides a set of results which are likely be close to optimal and potentially include the optimal solution. It should be also stressed that quantum annealers have a particular topology, meaning that a general QUBO can only be solved on them after having the respective Ising model’s adjacency graph embedded to the graph implemented by the hardware as an induced subgraph. This eventually requires multiple spins to be coupled to represent a singel spin (i.e. variable) of the problem.
@@
A physics-based alternative direction in solving QUBOs is the application of quantum-inspired algorithms [32]. Following this path, we apply the new VeloxQ solver [23], which appears to be effective for large, sparse QUBOs (with a limited number of connections) [26, 25, 27, 24]. This heuristic solver takes a QUBO or Ising instance and returns a series of solutions intended to originate from the problem’s low-energy spectrum (hopefully including at least one feasible solution). VeloxQ is a physics-inspired classical solver for QUBO/Ising problems that mimics energy-minimization dynamics without requiring quantum hardware and achieves high scalability on standard CPUs/GPUs. Compared to D-Wave quantum annealing and simulated bifurcation, it delivers comparable or superior solution quality and runtime, especially for large problem instances. As the solver does not need embedding, is efficient for problems of arbitrary connectedness.
@@ 3.1 Toy model
Let us now demonstrate in detail the actual behavior of our model on a toy example that can be followed in detail. The toy model is defined by the timetable $T=\{\tau_{1},\tau_{2},\tau_{3}\}$ where $\tau_{1}$ and $\tau_{2}$ are trains that go $A\rightarrow B$, and $\tau_{3}$ goes $B\rightarrow A$. The timetable is arranged so that any rolling stock that arrives at $B$ can either remain there or return to $A$. Two EMU types $R=\{r_{1},r_{2}\}$ are available at the depot at station $A$ throughout the time horizon, with fleet sizes $N(v_{d})_{r_{1}}=2,N(v_{d})_{r_{2}}=1$ with the seats capacity $p_{1}=70$, $p_{2}=110$. There is also an optional service train $v_{4}$ going from $B$ to $A$ (which can be served only by type $r_{1}$). We assume that $A=2$ drivers are available throughout the horizon; they can drive any of the two EMUs and take any of the trips. Finally, we do not consider bicycles in this example.
@@ 3.2 Practical problems
Let us discuss the results obtained by the VeloxQ quantum-inspired solver [23]. We observe better performance than D-Wave for small instances; however, large instances were not tractable with the QUBO approach because of the explosion of the number of terms. In fact it is a general observation of with the QUBO formulation: regardless of the QUBO solver applied, the main limitation on a bigger scale is the generation and storage of the QUBO model itself. Even the generation of the model can take more time than the ILP solution.
In those cases, however, when the QUBOs could be reasonably generated, VeloxQ gave promising results, outperforming DWave. VeloxQ has two parameters: the number of time steps and the number of repetitions, which we have set to the default $4096$ and $5000$, respectively, resulting in a compute time of $0.2$ seconds per instance in the cloud. Within this time, it has found the optimal solutions (as deduced from the exact ILP solutions) in all the cases which were tractable with the QUBO approach. This illustrates that a good QUBO heurisitc such as VeloxQ can be competitive even with ILP approaches, however, in the present problem the overhead in the problem generation poes a severe limitation: from instance $3$ on, the generation of the QUBOs was not reasonable.
@@ 4 Discussion
As of VeloxQ, a physics-inspired, cloud-based heuristic solver, in those cases when the generation of the QUBO could be performed with reasonable resources, the optimal solution, and a number of feasible alternatives close to it were found in a short and predefined computational time, illustrating that QUBO solvers themselves can be competitive, but they should be applied to the right problem. 4095 A key requirement from practice is not only to obtain one feasible plan, but to have multiple feasible alternatives available quickly so that dispatchers can account for non-modelled considerations and coordinate with infrastructure management. The toy example illustrates that the model naturally admits multiple feasible “excited” solutions beyond the optimum. This feature aligns well with the behavior of annealing-based approaches, which aim to sample from low-energy regions rather than output a single deterministic solution; in principle, this could be operationally valuable when a DSS needs to offer a small portfolio of near-optimal circulation options (e.g., different reserve usage levels, different coupling choices, or different depot return patterns) under time pressure. However, the experiments also show that obtaining a useful portfolio of feasible solutions from QUBO samplers requires careful constraint encoding (penalty design), feasibility checks, and likely post-processing—otherwise low-energy samples can still violate critical constraints (e.g., slack-variable feasibility).